OpenStudy (anonymous):

Can someone help explain why the limit of (10^n)/((-9)^(n-1)) is infinity as n goes towards positive infinity?

OpenStudy (johnweldon1993):

Hmm...well what I would do...is notice that \[\large \frac{10^n}{9^{n -1}}\] You can always write the denominator as a negative exponent to make this multiplication so that would become \[\large 10^n \times 9^{-(n-1)}\] Distribute that '-' sign... \[\large 10^n \times 9^{(1 - n)}\] And the plug in the infinity....well a really big number times a really big number is still a really big number...so we arrive at infinity again